if alpha and beta are the zeros of a quadratic polynomial 3x2+5x+7 , then 1/alpha cu. + 1/beta cu. is ?

Grade:Select Grade

2 Answers

Suresh Godi
8 years ago
1/a3 + 1/b3 = (a3 + b3 )/(ab)3 = [(a+b)3 -3ab(a+b)] / (ab) .
 
ab = product of the roots
a+b = sum of the roots
Suresh Godi
8 years ago
1/a3 + 1/b3 = (a3 + b3 )/(ab)3 = [(a+b)3 -3ab(a+b)] / (ab) .
 
ab = product of the roots
a+b = sum of the roots